OpenStudy (anonymous):

Maple syrup worth $6.00 a gallon and corn syrup worth .80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are need to make 50 gallons of the mixture?????

OpenStudy (anonymous):

X+Y=50 (gallons) 6.00X + 0.8Y = 118 (because 50* 2.36= 118$) therefor X+Y-50=0 and 6X+0.8Y-118=0 also 7X + 1.8Y - 168=0 (combining the two) 15X+1.8Y- 295= 0 (equation 2 times 2.5) (then subtract from other) 8X-127=0 X=15.875 Y= 50-X= 34.125 So you need 15.875 gallons of Maple and 34.125 gallons of Corn.

OpenStudy (anonymous):

that is total bull. let X represent the decimal percentage of corny syrup in the mix \[(0.00 \le X \le 1.00)\] so, for 1 gallon, 0.80X + 6.00(1.00-X) = 2.36 0.80X + 6.00 - 6.00X = 2.36 6.00 - 5.20X = 2.36 3.64 = 5.20X X = 3.64/5.20 X = 0.70 Therefore, for 50 gallons of the mix, 50*0.70 gallons of corn syrup are used and 50*(1-0.70) of maple are used. So, exactly 35 gallons of corn syrup are needed and 15 gallons of maple syrup are needed.